# Deep-learning-based continuous attacks on QKD protocols

### v2

@@ Appendix A Analytical derivation of the QBER without attacks
Appendix B Analytical derivation of Eve accuracy for a projective measurement
The result above allows us to easily determine Eve accuracy in the case she performs a projective measurement at time $t^{*}$. Indeed, when the photon reaches Eve at time $t^{*}$, she also chooses randomly one of the two available bases. The probabilities above correspond to someone measuring in the right basis. However, if Eve does not, her probability of detecting the right state is still $1/2$ as she can random guess. Thus, the probability that Eve deduces the right state (i.e., the accuracy $A$) can be obtained from Eq. (24), which yields
@@ Appendix B Analytical derivation of Eve accuracy for a projective measurement
Appendix C Analytical derivation of Bob accuracy for a projective measurement (intercept-and-resend attack)
@@
Appendix D Confidence of the model for $\sigma_{z}$ and $\sigma_{x}$ measurements
@@
Appendix E LSTM hidden states activation through time
For the initial state |$\ket{0}$ (Fig. 10(a)), some hidden states of the LSTM network activate at the early times of the time-series before maintaining a non-zero activity during the whole time sequence. However, for the initial state $\ket{+}$ (Fig. 10(b)), the hidden states retain their activity until the end of the time series, which could indicate that more information about the initial state is contained in the later times of the photocurrents. The early activation of the hidden states in both cases demonstrates that most information is contained in the first time steps of the currents, and is memorized by the hidden units until the output.

### v3

@@ Appendix B Analytical derivation of the QBER without attacks
Appendix C Analytical derivation of Eve accuracy for a projective measurement
The result above allows us to easily determine Eve accuracy in the case she performs a projective measurement at time $t^{*}$. Indeed, when the photon reaches Eve at time $t^{*}$, she also chooses randomly one of the two available bases. The probabilities above correspond to someone measuring in the right basis. However, if Eve does not, her probability of detecting the right state is still $1/2$ is still 1/2 as she can get each result with equal probability. Thus, the probability that Eve deduces the right state (i.e., the accuracy $A$) can be obtained from Eq. (47), which yields
@@ Appendix C Analytical derivation of Eve accuracy for a projective measurement
Appendix D Analytical derivation of Bob accuracy for a projective measurement (intercept-and-resend attack)
@@
Appendix E Confidence of the model for $\sigma_{z}$ and $\sigma_{x}$ measurements
@@
Appendix F LSTM hidden states activation through time
For the initial state |$\ket{0}$ (panel (a) of Fig. 10), almost all hidden states of the LSTM network activate at the early times of the time-series before maintaining a non-zero activity during the whole time sequence. However, for the initial state $\ket{+}$ (panel (b) of Fig. 10), the hidden states that activate retain their activity until the end of the time series, which could indicate that more information about the initial state is contained in the later times of the photo currents. The early activation of the hidden states in both cases demonstrates that most information is contained in the first time steps of the currents, and is memorized by the hidden units until the output.
@@ Appendix F LSTM hidden states activation through time
Appendix G Information gain computation
We start from the general definition of expected mutual information (i.e., information gain) for two random variables $X$ and $Y$ [73, 97]:
[…]
Let $S$ be a discrete variable representing the different initial states ($s=0,1,2,3$) and $X\in\mathbb{R}^{40}$ a $40$-dimensional continuous variable representing the values of the homodyne currents at each time step. The joint distribution can be written $P_{X,S}(x,s)=P(s)P(x|s)$, such that Eq. (55) becomes
[…]
It can also be expressed as the expected Kullback-Leibler (KL) divergence [97]:
$I(S;X)=\mathbb{E}_{s\sim P(s)}\left[D_{\mathrm{KL}}(P(x|s)\,\|\,P(x))\right].$ (57)
To compute Eq. (56), we assume that for each state $s$, the measurement vectors follow a multivariate Gaussian distribution
$P(x\mid s)=\mathcal{N}\left(x\mid\mu_{s},\Sigma_{s}\right).$ (58)
We test this assumption with the Henze-Zirkler multivariate normality test (over $2\times 10^{4}$ samples), which tests the null hypothesis that the data follows a multivariate normal distribution [98]. The results are displayed in Table 3.
State p-value
|$\ket{0}$ $0.8757$
|$\ket{1}$ $0.7008$
$\ket{+}$ $0.9424$
$\ket{-}$ $0.7862$
Table 3: P-values for the Henze-Zirkler multivariate normality test. A p-value above $0.05$ suggests the data could plausibly follow a multivariate normal distribution.
Conditional distributions are shown for each state in Fig. 11.
[…]
The logarithms in these formulae are usually taken to base 2 if information is measured in units of bits, or to base e if information is measured in nats. The resulting information gain is
$I(S;x)$ $=0.2894\ \text{nats}=0.4175\ \text{bits}.$